# Explicitly Modeling Subcortical Vision with a Neuro-Inspired Front-End Improves CNN Robustness

### v1

@@ Appendix A VOneNets
VOneNets [Dapello2020.06.16.154542] are convolutional neural networks (CNNs) augmented with a biologically-inspired, fixed-weight front-end simulating primary visual cortex (V1), termed the VOneBlock. This front-end is structured as a linear–nonlinear–Poisson (LNP) cascade [RUST2005945], incorporating a Gabor filter bank (GFB) [Jones1987-dy], nonlinearities for both simple and complex cells [Adelson:85], and a stochastic spiking mechanism modeling neuronal variability [Softky1993-te] (cf. Fig. A1). The GFB parameters are sampled from empirical distributions of orientation preference, spatial frequency (SF) tuning, and receptive field (RF) size [DEVALOIS1982545, DEVALOIS1982531, ringach2002spatial_dtructure]. The channels are split evenly between simple and complex cells, and a Poisson-like noise generator is applied to emulate spiking variability. The full implementation is available at https://github.com/dicarlolab/vonenet under a GNU General Public License v3.0.
@@
Figure A1: VOneNets simulate V1 processing upstream of standard CNNs. Each VOneNet incorporates a biologically-constrained front-end, the VOneBlock, preceding a conventional CNN. The VOneBlock consists of a fixed-weight Gabor filter bank (GFB) parameterized by empirical distributions, nonlinearities emulating simple and complex cell responses, and a stochastic component that injects Poisson-like noise to mimic V1 neuronal variability. Adapted from Baidya et al. [baidya2021combining].
@@
To construct the VOneNet variants, we replaced the initial block of each base architecture with the VOneBlock, along with a channel-matching bottleneck layer to maintain architectural compatibility between the front-end and the downstream convolutional stack. Consistent with the original VOneResNet50 model [Dapello2020.06.16.154542], we replaced a single convolutional layer, batch normalization, nonlinearity, and a max-pooling operation. When using an EfficientNet-B0, we substituted the initial convolution, batch normalization, and activation with the VOneBlock. For VOneResNet50, we preserved the original configuration of 512 channels within the VOneBlock, allocating 256 channels to each cell type (simple and complex). However, given the reduced channel dimensionality in the EfficientNet-B0 where the second stage expects only 32 channels, in contrast to 64 in ResNet50 we downscaled the VOneBlock accordingly, employing 128 channels per cell type.
@@
We modified the VOneBlock by adjusting the field of view (FoV) to 7deg (down from the original 8deg) and increased the spatial frequency range of the GFB to 0.5 – 8.0 cpd (from 0.5 – 5.6cpd). Additionally, to maintain consistency with upstream processing, we configured the GFB to uniformly sample a single channel from the input, regardless of any preceding subcortical transformations. Finally, to ensure methodological consistency in the spike-based activation regime across both the SubcorticalBlock (cf. Section E.2) and the VOneBlock, we imposed a unified temporal integration window of 50ms. In alignment with Table C2 of Dapello et al. [Dapello2020.06.16.154542], we applied a linear scaling factor to the VOneBlock outputs such that the mean evoked response to a batch of natural images from ImageNet matched the target stimulus response of 0.655 spikes. This variant of the VOneBlock was used on the both the VOneNet and the EVnet models described in this study.
@@ Appendix B Primate Vision Alignment
Shape-bias.
@@
In contrast to humans, who predominantly rely on shape cues for object recognition, ImageNet-trained CNNs have been shown to exhibit a strong bias toward texture-based representations [geirhos2022imagenettrained]. Measuring shape bias thus serves as a proxy for alignment with human inductive biases. We evaluate this using the cue conflict dataset from Geirhos et al. [geirhos2022imagenettrained], where images contain conflicting shape and texture cues (e.g., a cat-shaped image with elephant texture). While humans tend to classify by shape, ImageNet-trained CNNs often prefer texture. A model’s shape bias is computed as the proportion of shape-consistent predictions out of all shape- or texture-consistent responses.
BrainScore.
@@
The BrainScore platform [SchrimpfKubilius2018BrainScore, Schrimpf2020integrative] is a standardized benchmarking suite for evaluating how brain-like artificial neural networks (ANNs) are. In the context of object recognition, BrainScore compares model activations against neural recordings from primate visual areas and human behavioral data. For early visual processing, V1 predictivity is quantified via the FreemanZiemba2013 [Freeman2013] neural benchmark, while response properties in V1 are assessed using the Marques2020 [Marques2021.03.01.433495] benchmark. BrainScore aggregates multiple such benchmarks into a composite score that reflects a model’s alignment with neural and behavioral patterns observed in biological systems. Researchers can submit their models for evaluation at https://www.brain-score.org/.
V1 Predictivity.
To predict model’s ability to predict single-neuron responses in V1, we employed a dataset [Freeman2013] comprising responses from 102 V1 neurons to 450 unique 4deg image patches, spanning both naturalistic textures and noise stimuli. Predictivity was measured as the explained variance using partial least squares (PLS) regression under a 10-fold cross-validation scheme
V1 Response Properties.
Marques et al. [Marques2021.03.01.433495] introduced a novel model-to-brain comparison framework that bypasses conventional fitting procedures, instead relying on in silico neurophysiology to establish direct, one-to-one correspondences between artificial and V1 neurons. By probing models with canonical stimulus sets such as drifting gratings and texture pattern, the method quantifies alignment through a normalized similarity metric grounded in the Kolmogorov-Smirnov distance, capturing the distributional match of neural response properties. Critically, this framework enables rigorous benchmarking against prior neurophysiological studies without requiring raw recordings, effectively transforming existing literature into executable V1-aligning tests. In total, this method focuses on 22 distinct response characteristics, organized into seven functional domains: orientation tuning, spatial frequency tuning, receptive field size, surround modulation, texture modulation, response selectivity, and response magnitude. Table B1 presents all individual response properties along with the scores obtained for the ResNet50, the VOneResNet50 and EVResNet50 models.
@@
Table B1: EVResNet50 overperforms baselines on V1 response property alignment. BrainScore [SchrimpfKubilius2018BrainScore, Schrimpf2020integrative] V1 alignment scores for ResNet50, VOneResNet50, and EVResNet50. Values indicate mean $\pm$ SD ($n=3$ seeds).
@@
Category Resp. Property ResNet50 [he2015deep] VOneResNet50 EVResNet50
Orientation Orientation Selective 0.975$\pm$0.024 0.999$\pm$0.001 0.994$\pm$0.003
Circ. Variance (CV) 0.818$\pm$0.011 0.742$\pm$0.013 0.807$\pm$0.009
Orth./Pref. Ratio 0.855$\pm$0.023 0.717$\pm$0.014 0.755$\pm$0.008
CV Bandwidth Ratio 0.740$\pm$0.024 0.763$\pm$0.005 0.813$\pm$0.007
Pref. Orientation 0.943$\pm$0.046 0.985$\pm$0.004 0.976$\pm$0.015
Orth./Pref.-CV Diff. 0.766$\pm$0.016 0.885$\pm$0.004 0.857$\pm$0.008
Or. Bandwidth 0.659$\pm$0.086 0.922$\pm$0.010 0.964$\pm$0.006
Spatial Peak SF 0.551$\pm$0.047 0.961$\pm$0.002 0.957$\pm$0.005
Frequency SF Bandwidth 0.826$\pm$0.019 0.962$\pm$0.006 0.899$\pm$0.007
SF Selective 0.886$\pm$0.053 0.983$\pm$0.005 0.881$\pm$0.007
Response Texture Selective 0.678$\pm$0.008 0.800$\pm$0.004 0.821$\pm$0.006
Selectivity Modulation Ratio 0.349$\pm$0.009 0.737$\pm$0.002 0.875$\pm$0.005
Texture Var. Ratio 0.794$\pm$0.014 0.703$\pm$0.011 0.720$\pm$0.006
Texture Sparseness 0.663$\pm$0.032 0.927$\pm$0.002 0.939$\pm$0.003
RF Size Grating Sum. Field 0.272$\pm$0.005 0.547$\pm$0.016 0.705$\pm$0.009
Surround Diameter 0.156$\pm$0.000 0.361$\pm$0.015 0.733$\pm$0.003
Surround Mod. Surround Sup. Index 0.389$\pm$0.023 0.373$\pm$0.003 0.635$\pm$0.007
Texture Abs. Texture Mod. Idx. 0.978$\pm$0.019 0.942$\pm$0.004 0.845$\pm$0.009
Modulation Texture Mod. Idx. 0.606$\pm$0.040 0.897$\pm$0.011 0.627$\pm$0.010
Response Max. Texture 0.939$\pm$0.002 0.906$\pm$0.010 0.948$\pm$0.005
Magnitude Max. DC 0.873$\pm$0.053 0.824$\pm$0.008 0.902$\pm$0.023
Max. Noise 0.783$\pm$0.018 0.923$\pm$0.006 0.949$\pm$0.011
@@ C.1 Adversarial Attacks
To evaluate white-box robustness, we employed Projected Gradient Descent (PGD) [madry2019deep] on top of a subset of 5000 images from the ImageNet validation split. PGD is a widely adopted first-order attack that has proven effective against biologically inspired models such as VOneNets [Dapello2020.06.16.154542]. We selected this attack due to its scalability to large datasets (e.g., 5k ImageNet samples), and its compatibility with deterministic inference pipelines, which avoids the pitfalls of stochastic defenses that may artificially degrade attack success rates [Dapello2020.06.16.154542]. We run PGD for 16 iterations, where each step follows the update rule:
@@
where $\mathbf{x}^{t}$ denotes the adversarial input at iteration $t$, $\mathcal{L}$ is the cross-entropy loss function, and $\mathcal{P}_{\mathbf{x}+\mathcal{S}}$ projects back onto the perturbation set $\mathcal{S}$ centered at the clean input $\mathbf{x}$. Under an $L_{\infty}$ threat model, $\mathcal{S}$ corresponds to a box constraint, while for $L_{1}$ or $L_{2}$ norms with a perturbation budget $\epsilon$, the gradient direction is rescaled at each iteration to have the respective norm $\alpha$, and the projection ensures the final adversarial input $\mathbf{x}_{\text{adv}}$ satisfies $\|\mathbf{x}_{\text{adv}}-\mathbf{x}\|_{p}\leq\epsilon$. We used a used a step size of $\alpha=\|\mathbf{x}_{\text{adv}}-\mathbf{x}\|_{p}/8$ and performed a total of 12 attacks carried out under $L_{\infty}$, $L_{2}$ and $L_{1}$ norm constraints at four perturbation budgets each: $\|\delta\|_{\infty}\in[1/1020,1/255,4/255,16/255]$, $\|\delta\|_{2}\in[0.15,0.6,1.2,2.4]$ and $\|\delta\|_{1}\in[40,160,640,2560]$. We used the Adversarial Robustness ToolBox v1.17.1 [art2018] to conduct all the attacks.
@@
Due to the inherent stochasticity in our models, special considerations were necessary to enable gradient-based adversarial optimization. We first applied the reparameterization trick [Kingma2014], which permits gradient flow through stochastic nodes by expressing random variables as deterministic functions of noise. To obtain reliable gradient estimates for PGD, we further adopted the approach of Athalye et al. [obfuscated-gradients], replacing $\nabla f$ with an average over multiple stochastic forward passes. Specifically, we estimate gradients as
$\nabla f\approx\frac{1}{N}\sum_{i=1}^{N}\nabla_{i}f,$
@@
where each $\nabla_{i}f$ corresponds to a gradient computed using an independent Monte Carlo sample. We set $N=32$, increasing the number of samples compared to prior work on VOneNets (which used $N=10$) to compensate for the additional sources of noise introduced in our architecture.
@@
Figure C1: Increasing the number of Monte Carlo gradient samples (k) has limited impact on white-box attack effectiveness. White-box PGD accuracy ($\|\delta\|_{\infty}=1/255$, 16 iterations) evaluated on 500 images. Increasing $k$ from 32 to 128 leads to only marginal decreases in accuracy for both VOneResNet50 and EVResNet50, indicating that additional samples do not substantially strengthen the attack. Lines indicate the mean accuracy and shaded regions denote SD ($n=3$ seeds).
@@
To verify the reliability of our adversarial evaluation pipeline, we conducted a suite of sanity checks on a 500-image subset from the ImageNet validation set. In line with the recommendations of Athalye et al. [obfuscated-gradients] and Carlini et al. [carlini2019evaluating], we confirmed that top-1 accuracy decreases monotonically as a function of perturbation strength across all norm constraints (Tab. 4). Additionally, we also verified that increasing the number of gradient samples in the Monte Carlo approximation did not lead to a substantial increase in attack success (Fig. C1), further supporting the completeness of our threat model.
@@ C.2 Image Corruptions
Figure C3: EVResNet50 enhances robustness to noise, blur, and digital corruptions. Top-1 accuracy across corruption severity levels for the four categories of common corruptions: noise, blur, weather, and digital. Solid lines indicate the mean performance; error bars denote SD ($n=3$seeds)
The ImageNet-C dataset [hendrycks2019benchmarking] consists of 15 different corruption types, each at 5 levels of severity for a total of 75 different perturbations applied to validation images of the ImageNet validation split. Accuracy improvement on these datasets should be indicative of model robustness gains, given that it comprises, in total, 75 diverse corruptions. The individual corruption types are Gaussian noise, shot noise, impulse noise, defocus blur, glass blur, motion blur, zoom blur, snow, frost, fog, brightness, contrast, elastic transform, pixelated, and JPEG compression. The individual corruption types are grouped into 4 categories: noise, blur, weather, and digital effects. Examples of image corruptions are presented in Figure C2. The ImageNet-C dataset is publicly available at https://github.com/hendrycks/robustness under Creative Commons Attribution 4.0 International.

### v2

@@ Appendix A VOneNets
VOneNets [12] are convolutional neural networks (CNNs) augmented with a biologically-inspired, fixed-weight front-end simulating primary visual cortex (V1), termed the VOneBlock. This front-end is structured as a linear–nonlinear–Poisson (LNP) cascade [74], incorporating a Gabor filter bank (GFB) [75], nonlinearities for both simple and complex cells [76], and a stochastic spiking mechanism modeling neuronal variability [55] (cf. Fig. A1). The GFB parameters are sampled from empirical distributions of orientation preference, spatial frequency (SF) tuning, and receptive field (RF) size [47, 77, 78]. The channels are split evenly between simple and complex cells, and a Poisson-like noise generator is applied to emulate spiking variability. The full implementation is available at https://github.com/dicarlolab/vonenet under a GNU General Public License v3.0.
@@
Figure A1: VOneNets simulate V1 processing upstream of standard CNNs. Each VOneNet incorporates a biologically-constrained front-end, the VOneBlock, preceding a conventional CNN. The VOneBlock consists of a fixed-weight Gabor filter bank (GFB) parameterized by empirical distributions, nonlinearities emulating simple and complex cell responses, and a stochastic component that injects Poisson-like noise to mimic V1 neuronal variability. Adapted from Baidya et al. [79].
@@
To construct a VOneNet, we replaced the initial block of each base architecture with the VOneBlock, along with a channel-matching bottleneck layer to maintain architectural compatibility between the front-end and the downstream convolutional stack. Consistent with the original VOneResNet50 model [12], we replaced a single convolutional layer, batch normalization, nonlinearity, and a max-pooling operation and preserved the original configuration of 512 channels within the VOneBlock, allocating 256 channels to each cell type (simple and complex). When using an EfficientNet-B0, we substituted the initial convolution, batch normalization, and activation with the VOneBlock. Since this initial block has a total stride of 2, we decreased the stride of the VOneBlock accordingly. Furthermore, given the reduced channel dimensionality in the EfficientNet-B0 where the second stage expects only 32 channels, in contrast to 64 in ResNet50 we downscaled the VOneBlock, employing 128 channels per cell type. Finally, for the CORnet-Z we removed a single convolutional layer, nonlinearity, and max-pooling operation. Since this first block has the same combined stride and output dimension as the ResNet50, no additional modifications were necessary.
@@
We modified the VOneBlock by adjusting the field of view (FoV) to 7deg (down from the original 8deg) and increased the SF range of the GFB to 0.5 – 8.0 cpd (from 0.5 – 5.6cpd). This modification allows us to better match empirical V1 distributions, while maintaining a similar safety margin with respect to the Nyquist SF. Additionally, to maintain consistency with upstream processing, we configured the GFB to uniformly sample a single channel from the input, regardless of any preceding subcortical transformations. Finally, to ensure methodological consistency in the spike-based activation regime across both the SubcorticalBlock (cf. Section E.2) and the VOneBlock, we imposed a unified temporal integration window of 50ms. In alignment with Table C2 of Dapello et al. [12], we applied a linear scaling factor to the VOneBlock outputs such that the mean evoked response to a batch of natural images from ImageNet matched the target stimulus response of 0.655 spikes. This scaling factor was computed independently for the VOneNet and for each EVnet variant described in this study.
@@ Appendix B Primate Vision Alignment
B.1 Empirical V1 Tuning Curves
Figure B2: Examples of empirical V1 tuning curves retrieved from the literature. SF, size, and contrast tuning curves (left to right) for example V1 neurons. Left SF tuning curve of a simple (gray) and complex (black) cell to drifting grating stimuli. Markers represent the total number of F1 responses to gratings of different SF normalized to the best response and the solid line depicts a quadratic fit for purposes of illustrating the tuning profile (adapted from Figure 10 in Schiller et al. [80]). Middle Size tuning curve of two complex cells of V1 with distinct degrees of surround suppression under high contrasts. Gray depicts the a cell form V1 layer 4B under 0.15 contrast and black represents a cell from V1 layer 6 under 0.31 contrast. Markers represent each cell’s F1 response to differently-sized gratings and the line depicts the predicted response of a fitted DoG model discussed in the original article (adapted from Figure 1 in Sceniak et al. [59]). Right Contrast tuning curve of two simple V1 cell from the least (gray) and most (black) contrast sensitive thirds of their respective population. Marks indicate F1 response and the solid line depicts a fitted response model discussed in the original article (adapted from Figure 2 in Sclar et al. [67]). Data points extracted via WebPlotDigitalizer [81].
B.2 Shape-bias
@@
In contrast to humans, who predominantly rely on shape cues for object recognition, ImageNet-trained CNNs have been shown to exhibit a strong bias toward texture-based representations [8]. Measuring shape bias thus serves as a proxy for alignment with human inductive biases. We evaluate this using the cue conflict dataset from Geirhos et al. [8], where images contain conflicting shape and texture cues (e.g., a cat-shaped image with elephant texture). While humans tend to classify by shape, ImageNet-trained CNNs often prefer texture. A model’s shape bias is computed as the proportion of shape-consistent predictions out of all shape- or texture-consistent responses.
B.3 BrainScore
@@
The BrainScore platform [42, 43] is a standardized benchmarking suite for evaluating how brain-like artificial neural networks (ANNs) are. In the context of object recognition, BrainScore compares model activations against neural recordings from primate visual areas and human behavioral data. For early visual processing, V1 predictivity is quantified via the FreemanZiemba2013 [41] neural benchmark, while response properties in V1 are assessed using the Marques2020 [40] benchmark. BrainScore aggregates multiple such benchmarks into a composite score that reflects a model’s alignment with neural and behavioral patterns observed in biological systems. Researchers can submit their models for evaluation at https://www.brain-score.org/.
@@
Table B1: Detailed results for V1 response property alignment. BrainScore [42, 43] V1 alignment scores for ResNet50, VOneResNet50, and EVResNet50. Values indicate mean $\pm$ SD ($n=3$ seeds).
@@
Category Resp. Property ResNet50 [4] VOneResNet50 EVResNet50
[…]
B.4 V1 Predictivity
To predict model’s ability to predict single-neuron responses in V1, we employed a dataset [41] comprising responses from 102 V1 neurons to 450 unique 4deg image patches, spanning both naturalistic textures and noise stimuli. Predictivity was measured as the explained variance using partial least squares (PLS) regression under a 10-fold cross-validation scheme
B.5 V1 Response Properties
Marques et al. [40] introduced a novel model-to-brain comparison framework that bypasses conventional fitting procedures, instead relying on in silico neurophysiology to establish direct, one-to-one correspondences between artificial and V1 neurons. By probing models with canonical stimulus sets such as drifting gratings and texture pattern, the method quantifies alignment through a normalized similarity metric grounded in the Kolmogorov-Smirnov distance, capturing the distributional match of neural response properties. Critically, this framework enables rigorous benchmarking against prior neurophysiological studies without requiring raw recordings, effectively transforming existing literature into executable V1-aligning tests. In total, this method focuses on 22 distinct response characteristics, organized into seven functional domains: orientation tuning, spatial frequency tuning, receptive field size, surround modulation, texture modulation, response selectivity, and response magnitude. Table B1 presents all individual response properties along with the scores obtained for the ResNet50, the VOneResNet50 and EVResNet50 models.
@@ C.1 Adversarial Attacks
To evaluate white-box robustness, we employed Projected Gradient Descent (PGD) [10] on top of a subset of 5000 images from the ImageNet validation split. PGD is a widely adopted first-order attack that has proven effective against biologically inspired models such as VOneNets [12]. We selected this attack due to its scalability to large datasets (e.g., 5k ImageNet samples), and its compatibility with deterministic inference pipelines, which avoids the pitfalls of stochastic defenses that may artificially degrade attack success rates [12]. We run PGD for $N=64$ iterations, where each step follows the update rule:
@@
where $\mathbf{x}^{t}$ denotes the adversarial input at iteration $t$, $\mathcal{L}$ is the cross-entropy loss function, and $\mathcal{P}_{\mathbf{x}+\mathcal{S}}$ projects back onto the perturbation set $\mathcal{S}$ centered at the clean input $\mathbf{x}$. Under an $L_{\infty}$ threat model, $\mathcal{S}$ corresponds to a box constraint, while for $L_{1}$ or $L_{2}$ norms with a perturbation budget $\epsilon$, the gradient direction is rescaled at each iteration to have the respective norm $\alpha$, and the projection ensures the final adversarial input $\mathbf{x}_{\text{adv}}$ satisfies $\|\mathbf{x}_{\text{adv}}-\mathbf{x}\|_{p}\leq\epsilon$. We used a used a step size of $\alpha=\epsilon/32$ and performed a total of 12 attacks carried out under $L_{\infty}$, $L_{2}$ and $L_{1}$ norm constraints at four perturbation budgets each: $\|\delta\|_{\infty}\in[1/1020,1/255,4/255,16/255]$, $\|\delta\|_{2}\in[0.15,0.6,1.2,2.4]$ and $\|\delta\|_{1}\in[40,160,640,2560]$. We used the Adversarial Robustness ToolBox v1.17.1 [82] to conduct all the attacks.
Figure C1: Adversarial robustness is evaluated at convergance of PGD iterations. Top-1 white-box accuracy iteration curves for PGD attacks with $\|\delta\|_{\infty}=1/255$, $\|\delta\|_{2}=0.6$, $\|\delta\|_{1}=160$ constraints for ResNet50, VOneResNet50 and EVResNet50 models, evaluated on 500 images. The step size was adjusted to be $\epsilon$ for 1 iterations, and $2\epsilon/N$, in the remaining cases. Increasing the number of PGD iteration steps increases attack effectiveness only up to roughly 32 iterations. Lines indicate the mean accuracy and shaded error bars denote SD ($n=3$ seeds).
@@
Due to the inherent stochasticity in our models, special considerations were necessary to enable gradient-based adversarial optimization. We first applied the reparameterization trick [64], which permits gradient flow through stochastic nodes by expressing random variables as deterministic functions of noise. To obtain reliable gradient estimates for PGD, we further adopted the approach of Athalye et al. [38], replacing $\nabla f$ with an average over multiple stochastic forward passes. Specifically, we estimate gradients as
$\nabla f\approx\frac{1}{k}\sum_{i=1}^{k}\nabla_{i}f,$
@@
where each $\nabla_{i}f$ corresponds to a gradient computed using an independent Monte Carlo sample. We set $k=10$, similarly to prior work on VOneNets, given that the additional noise source in the SubcorticalBlock did not mask the gradients any further.
@@
Figure C2: Increasing the number of Monte Carlo gradient samples has limited impact on white-box attack effectiveness. White-box PGD accuracy evaluated under $\|\delta\|_{\infty}=1/255$ with 64 PGD iterations on 500 images. Increasing $k$ from 10 to 64 leads to only marginal decreases in accuracy for both VOneResNet50 and EVResNet50, indicating that additional samples do not substantially strengthen the attack. Lines indicate the mean accuracy and error bars denote SD ($n=3$ seeds).
@@
To verify the reliability of our adversarial evaluation pipeline, we conducted a suite of sanity checks on a 500-image subset from the ImageNet validation set. In line with the recommendations of Athalye et al. [38] and Carlini et al. [39], we confirmed that top-1 accuracy decreases monotonically as a function of perturbation strength across all norm constraints (Tab. 4). Additionally, we verified that increasing the number of PGD iterations increased attack effectiveness (Fig. C1) and that increasing the number of gradient samples in the Monte Carlo approximation did not lead to a substantial increase in attack success (Fig. C2), further supporting the completeness of our threat model.
@@ C.2 Image Corruptions
Figure C4: Detailed results for common corruptions benchmarks. Top-1 accuracy across 5 severity levels for the 15 individual common corruptions of ImageNet-C. Lines indicate the mean top-1 accuracy and error bars denote SD ($n=3$ seeds)
The ImageNet-C dataset [6] consists of 15 different corruption types, each at 5 levels of severity for a total of 75 different perturbations applied to validation images of the ImageNet validation split. Accuracy improvement on these datasets should be indicative of model robustness gains, given that it comprises, in total, 75 diverse corruptions. The individual corruption types are Gaussian noise, shot noise, impulse noise, defocus blur, glass blur, motion blur, zoom blur, snow, frost, fog, brightness, contrast, elastic transform, pixelated, and JPEG compression. The individual corruption types are grouped into 4 categories: noise, blur, weather, and digital effects. Examples of image corruptions are presented in Figure C3. The ImageNet-C dataset is publicly available at https://github.com/hendrycks/robustness under Creative Commons Attribution 4.0 International.